Find the value of Z if the normal curve area between 0 and Z (positive) is 0.4332.
1
Expert's answer
2020-05-21T17:53:10-0400

We have that,

"0.4332 = \frac{1}{\sqrt{2\pi}}\intop_0^z e^{-\frac{z^2 }{2}} dz,"

the last integral is laplace function and all its values ​​are already calculated and you can see them here http://yunus.hacettepe.edu.tr/~ezgiyilmaz/files/tables.pdf

As we can see the value of our integral reached when z is equal to 1.5.


Answer Z = 1.5